A bag contain 7 yellow 3 red and 5 green marbles.Jack draws two marbles at random without replacement .What is the probability that one is red and other is green?
Answer by stanbon(75887)

A bag contain 7 yellow 3 red and 5 green marbles.Jack draws two marbles at random without replacement .What is the probability that one is red and other is green?
------
red/green patterns: 2
probability of each pattern: (3*5)/(15*14) = 1/14
--------------
Ans:: 2*(1/14) = 1/7
-----
Cheers,
Stan H.